23. The estimated cash flow for each alternative of mutually exclusives

alternative with MARR of 20% are as follows: Use the incremental

analysis of IRR method to make a recommendation.

Alternatives M1 M2 M3

Initial Cost $42,000 82,000 60,000

Annual Benefit $35,000 40,000 48,000

Annual operating cost $23,000 25,000 32,000

Salvage value $8,000 15,000 16,500

Useful Life 8 years 8 years 8 years

Rate of Return 25% 22.50% 23.80%


1
Expert's answer
2021-06-21T12:20:41-0400

For M1


"(42000-8000)\times(1+0.25)^8+8\times(35000-23000)=298656"

for M2


"(82000-15000)\times(1+0.225)^8+8\times(40000-25000)=459753"

for M3


"(60000-16500)\times(1+0.238)^8+8\times(48000-32000)=368024"

Answer: M2